# Intercept Cancer: Cancer Pre-Screening with Large Scale Healthcare Foundation Models

### v1

@@ 6.1 Scaling Laws of Pretraining on Healthcare Records
To pretrain effectively on healthcare records, we first conduct a thorough scaling law analysis of CATCH-FM. Specifically, we use IsoFLOP profiling [34]: pretraining models with various sizes at target FLOPs by varying the number of pretraining tokens.
@@
Scaling laws in healthcare foundation models resemble the scaling laws of large language models [34], highlighting the potential of large-scale EHR foundation models. While emergent capabilities need further study, we pretrain and finetune a series of compute-optimal CATCH-FM models across various FLOPs (Table 3) and evaluate their efficacy in cancer risk prediction. The benefits of compute-optimal pretraining are detailed in Appendix G.
| Model | Pancreatic (1201/30382, 3.95%) | Liver (1099/28353, 3.88%) | Lung (868/30153, 2.88%) | ||||||
| AUPRC | Speci. NPV. | Sensi. | AUPRC | Speci. NPV. | Sensi. | AUPRC | Speci. NPV. | Sensi. | |
| XGBoost | 68.8 | 99.6 97.8 | 46.6 | 69.2 | 99.6 97.9 | 47.1 | 64.5 | 99.7 98.3 | 42.3 |
| LightGBM | 72.9 | 99.6 98.0 | 51.3 | 73.1 | 99.6 97.8 | 52.3 | 68.9 | 99.7 98.4 | 47.0 |
| RETAIN | 11.6 | 80.7 97.3 | 46.0 | 8.8 | 85.4 97.4 | 44.3 | 2.1 | 100.0 97.1 | 0.0 |
| StageNet | 11.8 | 100.0 96.0 | 0.0 | 9.0 | 100.0 96.1 | 0.0 | 6.2 | 100.0 97.1 | 0.0 |
| BioGPT-347m | 55.0 | 98.8 97.7 | 44.1 | 55.4 | 99.3 97.5 | 38.4 | 58.5 | 99.4 98.3 | 44.5 |
| Qwen2.5-500m | 73.7 | 98.9 98.4 | 62.4 | 70.0 | 99.4 98.2 | 55.0 | 70.2 | 99.3 98.7 | 57.1 |
| CATCH-FM-160m | 81.2 | 99.2 98.7 | 70.7 | 76.6 | 99.3 98.5 | 63.4 | 71.5 | 99.6 98.6 | 54.7 |
| CATCH-FM-410m | 81.2 | 99.3 98.7 | 70.8 | 78.5 | 99.5 98.5 | 63.6 | 73.6 | 99.7 98.6 | 55.3 |
| CATCH-FM-720m | 82.0 | 99.2 98.8 | 73.1 | 76.5 | 99.3 98.6 | 66.1 | 69.1 | 99.2 98.8 | 61.9 |
| CATCH-FM-1b | 82.7 | 99.2 98.9 | 73.5 | 77.0 | 99.1 98.7 | 68.7 | 69.6 | 99.2 98.8 | 62.0 |
| CATCH-FM-2.4b | 82.5 | 99.2 98.9 | 73.7 | 76.9 | 99.1 98.7 | 69.0 | 69.6 | 99.3 98.9 | 63.3 |
@@ 6.2 Overall Results
Table 4 shows the overall performance of cancer risk prediction in NHIRD-Cancer. Cancer risk prediction is a challenging task. Many baselines fail to achieve meaningful performance. CATCH-FM is the only method that achieves 60%+ sensitivity on all three types of cancer while maintaining nearly 99% specificity and NPV, confirming its ability to identify high-risk patients for further screening (high sensitivity) and avoid unnecessary patient distress (high specificity).
@@
CATCH-FM outperforms XGBoost and LightGBM, strong tree models that often outperformed EHR foundation models [97], by 10+% sensitivity. It also significantly outperforms medical (BioGPT) and general (Qwen) large language models pretrained on the texts of NHIRD. Sect. 6.4 further studies the benefits of operating on medical codes. Note that the specificity-sensitivity trade-off can be easily adjusted for different scenarios by changing classification thresholds. Our studies show that CATCH-FM maintains its advantage across all thresholds. Additional details are found in Appendix F.
| Model | Pancreatic | Liver | Lung | ||||||
| AUPRC | Speci. NPV. | Sensi. | AUPRC | Speci. NPV. | Sensi. | AUPRC | Speci. NPV. | Sensi. | |
| Training Cohorts: all cancer patients, Testing Cohorts: first cancer patients | |||||||||
| Pos/Neg: | (559/30382 1.83%) | (528/28353 1.86%) | (419/30153 1.39%) | ||||||
| XGBoost | 59.2 | 99.8 98.8 | 38.6 | 65.2 | 99.8 98.9 | 44.1 | 61.8 | 99.9 99.1 | 41.8 |
| Qwen2.5-500m | 66.6 | 98.9 99.2 | 61.7 | 64.4 | 99.4 99.1 | 55.9 | 62.5 | 99.3 99.3 | 56.1 |
| CATCH-FM-160m | 56.8 | 99.3 99.1 | 52.5 | 48.9 | 99.3 98.9 | 43.9 | 40.7 | 99.6 99.1 | 34.9 |
| CATCH-FM-1b | 60.3 | 99.2 99.1 | 55.4 | 50.0 | 99.1 99.1 | 50.6 | 40.7 | 99.2 99.2 | 41.1 |
| CATCH-FM-2.4b | 60.0 | 99.2 99.2 | 57.6 | 51.4 | 99.1 99.1 | 51.7 | 41.4 | 99.3 99.2 | 45.5 |
| Training Cohorts: all cancer patients, Testing Cohorts: reoccurred cancer patients | |||||||||
| Pos/Neg: | (642/30382 2.11%) | (571/28353 2.01%) | (449/30153 1.49%) | ||||||
| XGBoost | 90.7 | 99.8 99.5 | 80.4 | 91.7 | 99.8 99.5 | 80.4 | 89.0 | 99.9 99.6 | 76.6 |
| Qwen2.5-500m | 67.2 | 98.9 99.2 | 63.4 | 61.4 | 99.4 99.1 | 54.0 | 63.7 | 99.3 99.3 | 53.2 |
| CATCH-FM-160m | 89.3 | 99.3 99.7 | 86.1 | 85.9 | 99.3 99.6 | 81.4 | 80.7 | 99.6 99.6 | 73.9 |
| CATCH-FM-1b | 91.4 | 99.2 99.8 | 89.3 | 86.9 | 99.1 99.7 | 85.1 | 80.5 | 99.3 99.7 | 79.0 |
| CATCH-FM-2.4b | 86.0 | 99.2 99.7 | 87.7 | 84.9 | 99.1 99.7 | 85.0 | 78.8 | 99.3 99.7 | 80.0 |
| Training Cohorts:first cancer patients, Testing Cohorts: first cancer patients | |||||||||
| Pos/Neg: | (452/28058 1.59%) | (509/26148 1.95%) | (399/27396 1.45%) | ||||||
| XGBoost | 26.3 | 99.9 99.5 | 10.2 | 36.2 | 99.8 98.4 | 22.6 | 27.6 | 99.9 98.7 | 10.3 |
| Qwen2.5-500m | 22.0 | 99.3 98.6 | 17.9 | 32.4 | 99.0 98.5 | 29.2 | 15.9 | 99.1 98.8 | 18.4 |
| CATCH-FM-160m | 42.4 | 99.4 98.9 | 37.4 | 39.3 | 99.4 98.7 | 33.8 | 33.0 | 99.5 99.0 | 29.6 |
| CATCH-FM-1b | 57.2 | 99.6 99.2 | 48.7 | 49.6 | 99.4 98.9 | 44.4 | 47.5 | 99.4 99.2 | 45.6 |
| CATCH-FM-2.4b | 61.3 | 99.6 99.2 | 53.4 | 52.8 | 99.2 99.0 | 49.3 | 49.6 | 99.4 99.2 | 47.7 |
Table 5 shows CATCH-FM’s performances when finetuned with all cancer patients or only with the first cancer patients. Over half of the target patients fall into the reoccurred group, where CATCH-FM achieves 80% sensitivity, effectively leveraging previous cancer diagnostics as strong risk indicators. Detecting first cases is much more challenging. While other methods fail to balance the effectiveness of the two groups, CATCH-FM achieves strong effectiveness on both first and reoccurred patients when finetuned with the two groups together.
The benefit of scale is most evident in predicting the risks of the first cancer patients. CATCH-FM-2.4b outperforms CATCH-FM-160m by 5–10% sensitivity when finetuned on all cancer patients, and by more than 15% when trained exclusively in first-time cancer cases. Scaled-up foundation models perform better on harder tasks and with fewer finetuning labels. We further examine CATCH-FM’s generalization in the next experiment. Results with limited labels are in Appendix M.
@@ 6.3 Generalization Ability
Across Hospital Site. Figure 4 shows the K-shot results of CATCH-FM on the EHRSHOT official pancreatic cancer leaderboard, using K positive and K negative on-site examples for tuning. Despite significant shifts in population, healthcare systems, and noise in medical code mapping, CATCH-FM achieves state-of-the-art AUPRC across all shots on the EHRSHOT leaderboard with a handful of on-site examples, maintains the best AUROC for K64, and ranks second only at K=128 with a marginal difference (<3%). Moreover, CATCH-FM-160m, at a similar scale, outperforms CLMBR (140m) in AUPRC across all shots. Although its AUROC is lower, scaling to CATCH-FM-2.4b significantly improves few-shot performance, showing the benefits of scale. Notably, CLMBR is pretrained at a similar scale but with on-site data: 2.57M patient records from Stanford Medicine, the same as EHRSHOT, while CATCH-FM is challenged by distribution shifts and noisy medical code mappings. These results demonstrate CATCH-FM’s potential to transfer across different hospital sites.
Across Population. This experiment evaluates CATCH-FM on cohorts when cancer prescreening targets the general population with much lower cancer prevalence [71], and on randomly sampled controls. Figure 4 plots the performance with varying positive ratio and shows CATCH-FM’s robustness to different population distributions. For those interested, Appendix includes additional experiments on CATCH-FM’s robustness across other cohort variations (H), different history exclusion window (I), demographics (J), and its performance on other diagnostic support tasks (K).
| Variations | AUPRC | Specificity | Sensitivity |
@@
| CATCH-FM | 81.2 | 99.2 | 70.7 |
| demographic-only features | 51.9 | 100 | 0 |
| No Pretrain | 71.2 | 99.3 | 55.7 |
@@
| w. Token Level Rel-Pos | 81.1 | 99.2 | 70.4 |
| w/o. time token | 80.0 | 99.2 | 70.3 |
@@
| demographic-only features | 52.9 | 100 | 0 |
| No Pretrain | 39.9 | 96.5 | 49.6 |
| Finetune Pythia | 59.4 | 98.7 | 51.8 |
| Pretrain from Scratch | 64.8 | 98.9 | 52.6 |
| Continual Pretrain Pythia | 64.6 | 98.2 | 57.3 |
6.4 Ablation Study
Table 6 shows ablation studies on different components of CATCH-FM. Modeling patient history is a key source of evidence for CATCH-FM. Demographic information alone is a poor indicator; pretraining is a key advantage of CATCH-FM; different model architectures, similar to observations in large language models, yield mild differences. More detailed ablations are presented in Appendix N
| Cancer | Features Learned by Sparse Autoencoder [27] with Description Generated by Node2Graph [25] |
@@ 6.4 Ablation Study
| Adenocarcinoma, Esophagomyotomy, Norvasc tablet, Isosorbide mononitrate |
6.5 Interpretability Analysis
This experiment leverages recently developed LLM interpretation method [27; 44] to understand the risk factors captured by CATCH-FM. We train a sparse autoencoder on the prediction layers of finetuned CATCH-FM and then leverage neuron-to-Graph (N2G) [25] to explain the top active latent features on positive cancer cases. Additional implementation details are found in Appendix L. Table 7 identifies features corresponding to top cancer risk factors captured by CATCH-FM. It includes not only trivial factors, such as Type 2 diabetes [20] and high blood pressure [77], but also non-trivial ones such as hypothyroidism, thyroid disorders, and Norvasc, with some indicators only presented at top medical journals less than ten years ago [69; 95], after our training data cutoff.
@@ 7 Conclusion
CATCH-FM provides a new cancer risk prediction methodology by pretraining and finetuning Transformers on millions of patients’ medical code sequences. Its effectiveness (60%+ sensitivity), low risk (99% specificity), and wide applicability (only requiring inference on EHR records) make it a natural fit for cancer pre-screening. It helps healthcare professionals efficiently decide whom and when to screen for cancer, potentially improving the effectiveness and coverage of cancer screening and ultimately, patient outcomes. Our experiments on NHIRD and EHRSHOT demonstrated the benefit of scale in pretraining EHR foundation models, their generalization ability across significantly different healthcare systems, and their ability to capture non-trivial cancer risk factors. We hope our findings, analyses, and open-source codes can inspire and facilitate further research and deployments in leveraging AI to solve real-world healthcare problems.
@@ Appendix A NHIRD Details
The Taiwanese National Health Insurance Research Database (NHIRD) is one of the largest and most comprehensive de-identified population-wide EHR datasets globally [35], covering over 99% of the Taiwanese population. It contains longitudinal medical records spanning over two decades, including patient demographics, diagnoses, prescriptions, clinical events, medical procedures, and hospital visits from all hospitals and medical facilities in Taiwan. Maintained by the National Health Insurance Administration in collaboration with the Ministry of Health and Welfare, the NHIRD comprises both registration files and hospital claim data submitted for reimbursement under the National Health Insurance (NHI) program.
@@
Data De-identification and Privacy. All NHIRD data are de-identified following strict protocols mandated by the National Health Insurance Administration (NHIA). Personally identifiable information (PII), such as names and ID numbers, is removed and irreversibly encrypted using non-public anonymization methods [53; 33]. Our study uses only anonymized demographic variables, e.g., age and gender, for model training. Given the level of anonymization and the coarse granularity of these attributes, the risk of patient re-identification is negligible. Hence, our use of NHIRD data fully complies with privacy standards and poses no ethical or legal concerns regarding patient confidentiality.
Correctness of NHIRD. The universal healthcare system in Taiwan enables frequent patient visits and interactions and generates rich, detailed, and longitudinal patient records. In 2023, NHIRD recorded over 380 million medical visits for 23 million individuals, averaging 16.5 visits per person [1]. These statistics, including consistent surgery and prescription rates, align closely with our dataset statistics in Table 1, confirming the correctness of our dataset. Most importantly, all data in NHIRD receive strict validation by the National Health Insurance Administration (NHIA), which enforces robust quality control measures to eliminate duplication, correct inconsistencies, and ensure patient-level accuracy. Additionally, coding standards are uniformly implemented across all healthcare providers nationwide, guaranteeing consistency and correctness at scale. Given this systematic validation, national standardization, and population-wide coverage, NHIRD is an exceptionally accurate and dependable data source trusted and used in research published at top peer-reviewed medical journals [93; 51; 89; 83].
Cancer Diagnosis Validity. Cancer diagnoses in the NHIRD are made by licensed physicians from all hospitals and medical facilities and undergo strict validation by the National Health Insurance Administration (NHIA) to ensure diagnostic accuracy, prevent misclassification, and eliminate billing errors or fraud [53; 35]. This rigorous quality control process makes the NHIRD a highly reliable source for cancer retrospective research, especially for creating cancer patient cohorts with diagnosis codes. With verified patient histories and clinically validated cancer diagnoses, the NHIRD has been widely adopted for cancer research and trusted in top peer-reviewed studies [54; 14; 37].

### v2

@@ 6.1 Scaling Laws of Pretraining on Healthcare Records
To pretrain effectively on healthcare records, we first conduct a thorough scaling law analysis of CATCH-FM. Specifically, we use IsoFLOP profiling (Hoffmann et al., 2022): pretraining models with various sizes at target FLOPs by varying the number of pretraining tokens.
@@
Scaling laws in healthcare foundation models resemble the scaling laws of large language models (Hoffmann et al., 2022), highlighting the potential of large-scale EHR foundation models. While emergent capabilities need further study, we pretrain and finetune a series of compute-optimal CATCH-FM models across various FLOPs (Table 3) and evaluate their efficacy in cancer risk prediction. The benefits of compute-optimal pretraining are detailed in Appendix F.
| Model | Pancreatic | Liver | Lung | ||||||
| first / subsequent (1.59% / 2.73%) | first / subsequent (1.95% / 2.25%) | first / subsequent (1.45% / 1.80%) | |||||||
| AUROC | AUPRC | Sensitivity | AUROC | AUPRC | Sensitivity | AUROC | AUPRC | Sensitivity | |
| XGBoost | 91.6 / 95.4 | 26.3 / 68.4 | 31.0 / 61.9 | 91.2 / 95.8 | 36.2 / 69.7 | 36.3 / 65.2 | 91.4 / 95.3 | 27.6 / 69.4 | 32.3 / 66.5 |
| LightGBM | 91.5 / 95.8 | 25.6 / 69.1 | 31.9 / 62.1 | 92.0 / 95.9 | 35.4 / 69.9 | 36.2 / 66.9 | 91.5 / 95.6 | 24.3 / 69.2 | 31.1 / 69.3 |
| RETAIN | 68.9 / 21.6 | 3.4 / 2.0 | 0.0 / 0.0 | 74.8 / 20.3 | 5.2 / 1.4 | 0.0 / 0.0 | 74.2 / 82.6 | 3.9 / 8.9 | 0.0 / 0.0 |
| StageNet | 64.3 / 69.7 | 2.4 / 4.5 | 0.0 / 0.0 | 59.9 / 64.1 | 2.4 / 3.1 | 0.0 / 0.0 | 66.9 / 64.1 | 2.3 / 3.1 | 0.0 / 0.0 |
| BioGPT-347m | 91.8 / 93.7 | 19.5 / 50.1 | 19.9 / 42.0 | 88.6 / 93.3 | 24.9 / 49.1 | 22.4 / 42.2 | 89.5 / 90.7 | 19.6 / 50.0 | 24.9 / 48.6 |
| Qwen2.5-500m | 90.3 / 92.7 | 22.3 / 57.9 | 25.4 / 50.8 | 90.4 / 93.7 | 32.4 / 60.3 | 32.4 / 55.9 | 86.3 / 92.8 | 15.9 / 60.1 | 18.8 / 53.1 |
[…]
@@ 6.2 Overall Results
Table 4 shows the overall performance of cancer risk prediction in NHIRD-Cancer. CATCH-FM outperforms XGBoost and LightGBM, strong tree models that often outperformed EHR foundation models Wornow et al. (2023), by 20%+ AUPRC on first target cancer cohorts and 15%+ AUPRC on subsequent target cancer cohorts. It also significantly outperforms medical (BioGPT) and general (Qwen) large language models pretrained on the texts of NHIRD. Sect. 6.4 further studies the benefits of pretraining directly on medical codes rather than converting codes into natural language. Moreover, our studies show that CATCH-FM achieve 50% and 70% sensitivity in first and subsequent cohorts at a decision threshold corresponding to 99% specificity, confirming its ability to identify high-risk patients for further screening (high sensitivity) and avoid unnecessary patient distress (high specificity). Additional details are found in Appendix E.
Cancer risk prediction is a challenging task, particularly for first target cancer cohorts, where many baselines fail to achieve meaningful performance. CATCH-FM demonstrates strong effectiveness, achieving 50%+ AUPRC in predicting first cancer cases. For subsequent target cancer cohorts, CATCH-FM effectively leverages prior cancer diagnoses as strong risk indicators and achieves 80%+ AUPRC. The benefit of scale is evident in all cohorts. CATCH-FM-2.4b outperforms CATCH-FM-160m by 10–15% AUPRC on first cancer cohorts and by 5–10% on subsequent cohorts. Scaled-up foundation models perform better on the harder first cancer prediction task.
| Methods | AUROC | AUPRC | Sensitivity |
| Cancer: Liver, Pos./Neg. (Ratio): 5798/297459 (1.95%) | |||
| XGBoost | 93.3 | 41.8 | 40.7 |
| Qwen2.5-500m | 91.9 | 30.0 | 29.1 |
| CATCH-FM-2.4b | 93.6 | 44.8 | 42.6 |
| Cancer: Lung, Pos./Neg. (Ratio): 5024/347974 (1.44%) | |||
| XGBoost | 93.7 | 34.7 | 37.9 |
| Qwen2.5-500m | 89.1 | 13.6 | 17.2 |
| CATCH-FM-2.4b | 94.1 | 36.4 | 41.4 |
@@ 6.3 Generalization Ability
We evaluate CATCH-FM under two generalization settings: temporal shifts within the same healthcare system (NHIRD), and distributional shifts across different systems and countries (EHRSHOT).
Across Hospital Site and Time. We evaluate CATCH-FM on cohorts from an adjacent timeline within the same hospital system, targeting first liver and lung cancer. Table 5 shows zero-shot results on NHIRD-Forward, demonstrating robustness to temporal shifts in population. Additional robustness analyses on cohort variations (G) and history exclusion windows (H) are in the Appendix.
Across Healthcare System and Country. Figure 4 shows the K-shot results of CATCH-FM on the EHRSHOT official pancreatic cancer leaderboard, using K positive and K negative on-site examples for tuning. Despite significant shifts in population, healthcare systems, and more than 50% medical code mismatch, CATCH-FM achieves state-of-the-art across all shots on the EHRSHOT leaderboard with a only handful of on-site examples, maintaining the best AUROC and AUPRC.
Moreover, CATCH-FM-160m outperforms CLMBR (140m) in AUROC for K 8 and in AUPRC, with only minor gaps at a lower shots, while scaling to CATCH-FM-2.4b further boosts few-shot performance, highlighting the benefits of scale. Notably, CLMBR is pretrained at a similar scale with on-site data (2.57M Stanford patients, same as EHRSHOT), while CATCH-FM faces distribution shifts with only 43% of SNOMED codes mapped to ICD. As ICD is already the predominant U.S. EHR standard (Feinstein et al., 2023), such SNOMED-to-ICD issues occur in only a fraction of healthcare providers. Overall, these results highlight CATCH-FM’s robustness to distribution shifts and code mapping loss, underscoring its potential for transfer across healthcare systems and sites.
6.4 Ablation Study
Table 6 shows ablation studies on different components of CATCH-FM. Modeling patient history is a key source of evidence for CATCH-FM. Demographic information alone is a poor indicator; pretraining is a key advantage of CATCH-FM; different model architectures, similar to observations in large language models, yield mild differences.
| Variations | AUROC | AUPRC | Sensitivity |
@@ 6.4 Ablation Study
| CATCH-FM | 91.4 | 42.4 | 43.1 |
| demographic-only features | 50 | 1.8 | 0.0 |
| No Pretrain | 86.9 | 16.7 | 18.7 |
@@
| w. Token Level Rel-Pos | 90.8 | 42.0 | 42.1 |
| w/o. time token | 90.6 | 41.2 | 41.3 |
@@
| demographic-only features | 50 | 2.0 | 0.0 |
| No Pretrain | 90.5 | 18.8 | 17.5 |
| Finetune Pythia | 89.9 | 16.7 | 18.8 |
| Pretrain from Scratch | 86.5 | 14.5 | 15.5 |
| Continual Pretrain Pythia | 91.4 | 25.8 | 29.4 |
| Cancer | Features Learned by Sparse Autoencoder Gao et al. (2024) with Description Generated by Node2Graph Foote et al. (2023) |
@@
| Esophagomyotomy, Norvasc tablet, Isosorbide mononitrate |
6.5 Risk Factors Captured
This experiment leverages recently developed LLM interpretation method (Gao et al., 2024; Kang et al., 2025) to understand the risk factors captured by CATCH-FM. We train a sparse autoencoder on the prediction layers of finetuned CATCH-FM on first target cancer patients and then leverage neuron-to-Graph (N2G) (Foote et al., 2023) to explain the top active latent features on positive cancer cases. Additional implementation details are found in Appendix I.
Table 7 identifies features corresponding to top cancer risk factors captured by CATCH-FM. It includes not only trivial factors, such as Type 2 diabetes (Cui & Andersen, 2012) and high blood pressure (Stocks et al., 2012), but also non-trivial ones such as hypothyroidism, thyroid disorders, and Norvasc, with some indicators only presented at top medical journals less than ten years ago (Sarosiek et al., 2016; Wang et al., 2018b), after our training data cutoff.
@@ 7 Conclusion
CATCH-FMprovides a new cancer risk prediction methodology by pretraining and finetuning Transformers on millions of patients’ medical code sequences. Its effectiveness (50%+ and 70%+ sensitivity on first and subsequent target cancer cohorts), low risk (99% specificity), and wide applicability (only requiring inference on EHR records) make it a natural fit for cancer pre-screening. It helps healthcare professionals efficiently decide whom and when to screen for cancer, potentially improving the effectiveness and coverage of cancer screening and ultimately, patient outcomes. Our experiments on NHIRD and EHRSHOT demonstrated the benefit of scale in pretraining EHR foundation models, their generalization ability across significantly different healthcare systems, and their ability to capture non-trivial cancer risk factors. We hope our findings, analyses, and open-source codes can inspire and facilitate further research and deployments in leveraging AI to solve real-world healthcare problems.
@@ Acknowledgments
Ethics Statement
Our research was conducted using de-identified EHRs obtained through formal research collaboration under regulated access and IRB review, with a determination of exemption made as the study relied solely on secondary, anonymized data and involved no human subjects. We ensured strict compliance with data governance rules set by the dataset sources for access. Privacy and confidentiality were preserved at all times, and no attempt was made to re-identify individuals. We restrict the release of model weights to authorized researchers with the same data access approvals, thereby safeguarding against misuse. Our work aims to advance early cancer detection through trustworthy artificial intelligence and foundation models while acknowledging potential biases in historical health data and the need for careful validation before any clinical deployment.
Reproducibility Statement
To ensure the reproducibility of our paper, a complete description of the NHIRD data schema, preprocessing pipeline, and full dataset statistics is detailed in the appendix. For transparency, we will publicly release our preprocessing, modeling, and training code under the MIT license, along with detailed instructions for reproducing the experiments. While direct redistribution of NHIRD data is restricted, all eligible institutions can apply for access under the same IRB framework. We will release pre-trained model checkpoints to authorized researchers with equivalent NHIRD data access for replication.
@@ Appendix A NHIRD Details
The Taiwanese National Health Insurance Research Database (NHIRD) is one of the largest and most comprehensive de-identified population-wide EHR datasets globally (Hsieh et al., 2019), covering over 99% of the Taiwanese population. It contains longitudinal medical records spanning over two decades, including patient demographics, diagnoses, prescriptions, clinical events, medical procedures, and hospital visits from all hospitals and medical facilities in Taiwan. Maintained by the National Health Insurance Administration in collaboration with the Ministry of Health and Welfare, the NHIRD comprises both registration files and hospital claim data submitted for reimbursement under the National Health Insurance (NHI) program.
@@
Data De-identification and Privacy. All NHIRD data are de-identified following strict protocols mandated by the National Health Insurance Administration (NHIA). Personally identifiable information (PII), such as names and ID numbers, is removed and irreversibly encrypted using non-public anonymization methods (Lin et al., 2018; Health & Center, ). Our study uses only anonymized demographic variables, e.g., age and gender, for model training. Given the level of anonymization and the coarse granularity of these attributes, the risk of patient re-identification is negligible. Hence, our use of NHIRD data fully complies with privacy standards and poses no ethical or legal concerns regarding patient confidentiality.
Correctness of NHIRD. The universal healthcare system in Taiwan enables frequent patient visits and interactions and generates rich, detailed, and longitudinal patient records. In 2023, NHIRD recorded over 380 million medical visits for 23 million individuals, averaging 16.5 visits per person (tai, 2024). These statistics, including consistent surgery and prescription rates, align closely with our dataset statistics in Table 1, confirming the correctness of our dataset. Most importantly, all data in NHIRD receive strict validation by the National Health Insurance Administration (NHIA), which enforces robust quality control measures to eliminate duplication, correct inconsistencies, and ensure patient-level accuracy. Additionally, coding standards are uniformly implemented across all healthcare providers nationwide, guaranteeing consistency and correctness at scale. Given this systematic validation, national standardization, and population-wide coverage, NHIRD is an exceptionally accurate and dependable data source trusted and used in research published at top peer-reviewed medical journals (Wang et al., 2018a; Lee et al., 2019; Tsai et al., 2024; Tain et al., 2025).
Cancer Diagnosis Validity. Cancer diagnoses in the NHIRD are made by licensed physicians from all hospitals and medical facilities and undergo strict validation by the National Health Insurance Administration (NHIA) to ensure diagnostic accuracy, prevent misclassification, and eliminate billing errors or fraud (Lin et al., 2018; Hsieh et al., 2019). This rigorous quality control process makes the NHIRD a highly reliable source for cancer retrospective research, especially for creating cancer patient cohorts with diagnosis codes. With verified patient histories and clinically validated cancer diagnoses, the NHIRD has been widely adopted for cancer research and trusted in top peer-reviewed studies (Lin et al., 2015; Chien et al., 2016; Huang et al., 2023).
@@
Subsequent Cancer Definition. In cancer registry standards such as SEER and IARC, the term subsequent primary cancer refers strictly to new independent primary malignancies, explicitly excluding metastases or recurrences of prior cancers. In contrast, for the purposes of this study, we define subsequent cancer more broadly to include both new independent primary cancers and metastases to the target organ. This operational definition reflects the screening context, where both scenarios represent clinically relevant risks for early detection.